OpenStudy (anonymous):

Pizza Problem: Medium (12 inches) Large (inches) Super (16 inchs) Price: Med: 8.99 Large: 11.99 Super: 13.99 1. Compare the area of each pizza to the nearest square inch. 2. Compute the amount of pizza per dollar for each pizza to the nearest tenth of an inch.

OpenStudy (anonymous):

The large pizza is 14 inches.

OpenStudy (turingtest):

what is the formula for the area of a circle?

OpenStudy (anonymous):

Area=pieR^2

OpenStudy (turingtest):

pie=tasty pastry pi=3.14159.... but yeah so using that formula, what is the area of each pizza?

OpenStudy (anonymous):

12, 14 and 16 inches right?

OpenStudy (anonymous):

No, sorry

OpenStudy (anonymous):

The area would be 113, 154 and 201 right?

OpenStudy (turingtest):

no first are the numbers given the radii of the pizzas, or the diameters?

OpenStudy (anonymous):

Diameter

OpenStudy (turingtest):

yeah those areas are right so find the differences between them

OpenStudy (anonymous):

0, 41, 47

OpenStudy (turingtest):

lol, the rest are right, but where did 0 come from ?

OpenStudy (turingtest):

the difference between the size of the small pizza and itself? haha

OpenStudy (anonymous):

Well the medium pizza is the base number?!

OpenStudy (turingtest):

I think they want like 201-154 154-113 201-113

OpenStudy (turingtest):

funny thought though :)

OpenStudy (anonymous):

So how do I calculate the amount of pizza per dollar?

OpenStudy (turingtest):

\[\frac{$}{\text{inches}^2}=\frac{\text{Total cost of that pizza}}{\text{Area of that pizza}}\]

OpenStudy (turingtest):

oh you wanted the other way 'round

OpenStudy (anonymous):

201-154=47 154-113=41 201-113=88

OpenStudy (turingtest):

\[\frac{\text{inches}^2}{$}=\frac{\text{Area of that pizza}}{\text{Total cost of that pizza}}\]

OpenStudy (turingtest):

yeah, looks earlier answer looks good

OpenStudy (anonymous):

113/$8.99=12.56

OpenStudy (turingtest):

yep units are square inches per dollar

OpenStudy (turingtest):

nice job

OpenStudy (anonymous):

So the right answer is 12.56 sq/in per dollar?

OpenStudy (turingtest):

\[12.56\frac{\text{in}^2}$\]or 12.56sqn/$ if yuo can't use latex

OpenStudy (anonymous):

12.56sqin/$, 12.84sqin/$, 14.36sqin/$

OpenStudy (anonymous):

Thanks again

OpenStudy (turingtest):

welcome :)